OpenStudy (anonymous):

Mixed Practice Solve the equation and check your solution 2/5+1/y-5=y-7/5y-25

OpenStudy (jmark):

Given 2/5+1/y-5=y-7/5y-25 then y has 2 value y=√2591/2-51/2 =-0.049061 y=-51/2-√2591/2 =-50.951 More information on Equation of a Sphere http://goo.gl/TvV26a